OpenStudy (anonymous):

Simplify 4 square root of 192

OpenStudy (anonymous):

does it help to know \(192=64\times 3\) ?

OpenStudy (dan815):

i fail at multiplication lol

OpenStudy (anonymous):

so \[\sqrt{192}=\sqrt{64\times 3}=\sqrt{64}\sqrt{3}=8\sqrt{3}\]

OpenStudy (jdoe0001):

so using satellite73's suggestion, your radical will look like \(\large 4\sqrt{64\times 3}\)

OpenStudy (dan815):

you can take it 1 step at a time too, if u cant see its 64*3 instantly like me..

OpenStudy (dan815):

first try taking it in half and so on, eventually ull get the right answer

OpenStudy (anonymous):

wait so the answer is not 8 sqr of 3?

OpenStudy (dan815):



OpenStudy (dan815):

thats the radical only

OpenStudy (jdoe0001):